# On the Soundness and Consistency of LLM Agents for Executing Test Cases Written in Natural Language

### v1

@@ 1 Introduction
We propose in the paper an algorithm and contexts to make NL test case unsoundness acceptable. We formulate this notion of acceptability and define measures to evaluate test case execution consistency. Specifically, we propose an algorithm for executing NL test cases with guardrail mechanisms. It dynamically completes the test steps and controls them, enabling evaluation of whether each step is executed as intended. We leverage a set of specialised agents to handle specific actions: navigation through the GUI, readiness of the GUI to perform the next step, and evaluation of assertions. To assess the capabilities of these agents in performing these actions, we define three measures. In addition, we introduce another measure to quantify the execution consistency of NL test cases—that is, the ability of our algorithm to produce stable and repeatable test outcomes.
@@
  • two prototype tools and corresponding test suites, made publicly available in [Salva and Taguelmimt, 2025].
@@ 3 Test Case Execution
With this in mind, we propose an algorithm that supplements on the fly a NL test case with additional internal actions. These actions aim at orchestrating the execution of agents and supervise how the $AUT$ behaves, ensuring that every navigation action is both feasible and carried out as expected. Specifically, we introduce two internal actions, denoted readinessreadiness and observeobserve, which are systematically injected for every navigation action, as illustrated in Figure 4(a). These actions are said to be internal in the sense that they are not observable during execution. readinessreadiness checks whether the next navigation action is executable, i.e., whether the necessary UI elements are present on the interface to perform the action. In turn, observeobserve checks whether the last executed navigation action has led to a visible change in the GUI and whether the agent has successfully completed it. Both internal actions also contribute to evaluating the consistency and reliability of the test case execution (Section 4). In addition to these internal actions, our algorithm adds to the test case the notion of inputs, i.e. actions starting by the symbol ”?” applied or sent to $AUT$, and outputs, i.e. observations initiated by $AUT$ the starting by the symbol ”!”.
All this is reflected in Algorithm 1. It incrementally unfolds the NL test case $tc$ into a tree structure expressing the synchronous execution of $tc$ on $AUT$. We propose to model this execution by using an Input Output Labelled Transition System (IOLTS) denoted $tc|AUT$ whose edges are labelled with actions and terminal nodes are labelled with test verdicts, $tc|AUT=<Q,L\cup\mathcal{T},\rightarrow,q_{0}>$, with $Q$ the state set, $L=L_{I}\cup L_{O}$ the set of inputs and outputs, respectively, $\mathcal{T}$ the set of internal actions, $\rightarrow$ the transition set and $q_{0}$ the initial state. Algorithm 1 starts by covering every navigation action $a_{i}$ in $a_{1}\dots a_{k}$. It encapsulates the action $a_{i}$ flanked by its corresponding actions $readiness(a_{i})$ and observeobserve, as illustrated in Figure 4(a). If either $readiness(a_{i})$ or observeobserve are evaluated to false, the test case cannot be executed because the action cannot be performed or has not updated the GUI, and an Inconclusive verdict is returned. Otherwise, the algorithm continues with the next action. When the algorithm reaches the assertions $\mathcal{A}_{k+1}\dots\mathcal{A}_{l}$, it incrementally evaluates them, as illustrated in Figure 4(b). If an assertion evaluates to false, then a Fail verdict is returned. The Inconclusive verdict is also reached in case an error is triggered by an agent. Otherwise, a Pass verdict is given.
@@ 3.1 Navigation Action Execution
The agent is invoked using prompts we wrote and structured following patterns proposed in [White et al., 2023]. We here considered: 1. the “Fact Checklist” pattern that asks the agent to output a list of facts included in the final answer, thereby enhancing transparency on the results provided; 2. the “Template” pattern that enforces an output format in JSON, which includes fields such as extracted facts, results, and a boolean indicator of task completion status; 3. The “Recipe” pattern, which builds a complete sequence of steps given some partially “ingredients” to achieve the given task. The prompt is given in Annex in Figure 8.
@@ 3.3 Assertion Evaluation
Table 2: Examples of assertions expressed with formulas.
@@ 4 Test Case Execution Consistency
We recall that a test case is made up of four kinds of actions: readinessreadiness, navigation, assertion, and observation. Among these, three are performed by agents—readinessreadiness, navigation actions, and assertions. To measure the consistency of their executions, we define three measures as follows:
First, $s_{r}(a)$ measures the consistency of $readiness(a)$, with $a$ a navigation action. Recall that the action $readiness(a)$ is either performed by readiness_strict(a)readiness\_strict(a) or readiness_w_agent(a)readiness\_w\_agent(a). readiness_strictreadiness\_strict is expressed by a formula and is consistent. When the formula evaluates to false, we use readiness_w_agent(a)readiness\_w\_agent(a), which itself calls $agent_{readiness}$. The execution consistency of readiness_w_agent(a)readiness\_w\_agent(a) is related to the standard deviation of $agent_{readiness}$. $s_{r}$ is hence defined as:
$0\leq s_{r}(a)=_{def}\left\{\begin{array}[]{l}1\text{ if }readiness\_strict(a)=1\\ 1-2\sigma(agent_{readiness})\text{ otherwise }\end{array}\right$.
@@
Third, $s_{a}(\mathcal{A})$ measures the consistency of the evaluation of $\mathcal{A})$. Similarly to readinessreadiness, this action is either performed by $assert\_strict(\mathcal{A})$ or $sert\_w\_agent(\mathcal{A}). sas_{a}$ is hence defined as:
$0\leq s_{a}(\mathcal{A})=\left\{\begin{array}[]{l}1\text{ if }assert\_strict(\mathcal{A})=1\\ 1-2\sigma(agent_{assert})\text{ otherwise }\end{array}\right$.
@@ 5 Test Case Unsoundness
Let $tc=a_{1}\dots a_{k}\mathcal{A}_{k+1}\dots\mathcal{A}_{l}$ be a NL test case such that there exists $a_{1}!g_{1}\dots?a_{k}!g_{k}\in traces(S)$, and $\mathcal{A}_{j}(k+1\leq j\leq l)$ are true on $g_{k}$, and $tc$ is made up of atomic actions and assertions. If $(\sigma(agent_{nav})<\sigma_{3}$, $\sigma(agent_{eval})<\sigma_{3}$, $\sigma(agent_{assert})<\sigma_{3})\in\Omega_{r}$, then $tc$ is weakly unsound.
@@
Proofs of these propositions are given in Annex.
@@ 6 Experimental Results
To perform these experiments, we developed 3 prototype tools composed of three agents. Two agents aim at evaluating readiness actions and assertions, as described in Section 3. The last agent was built on top of the Stagehand  [^3] https://www.stagehand.dev/ framework, which automates browser interactions with natural language and code by means of LLMs. Stagehand was used to perform navigation actions and, in part, for data extraction from pages:
@@
  • NLTestRunner implements Algorithm 1. Given a test suite and the number of runs $N$, it executes NL test cases, returns verdicts, and computes the estimated consistency measures and the observed test case execution consistencies if $N>0$. It can recognise 12 prompt forms to perform strict readiness and strict assertion evaluations.
We performed our experiments with 8 LLMs, whose sizes range between 3B and 70B parameters. The LLM inference was carried out using Ollama [^4] https://ollama.com/, installed on a local server equipped with an NVIDIA L40 GPU (48 GB memory). All Web sites, test suites, and tools are available in [Salva and Taguelmimt, 2025].
@@
Setup: To address this question, we considered 8 LLMs of varying sizes: Qwen2.5:3B, Qwen2.5 7B, Mistral 7B, Mistral Nemo 12B, Deepseek-R1 14B, Qwen3:14B, Mistral Devstral 24B, and Llama3.1:70B. These models, provided by different companies, range from 3B to 70B parameters. We selected LLMs that support tool calling, which is a required feature to allow GUI interaction.
@@
Results: Figures 5 and 6 present box-plots of the distributions of mean accuracies across the NL test cases. The plots also show quartiles and mean values (indicated with crosses) of performing navigation actions, returning expected readiness, and returning expected assertion verdicts. Table 3 lists the standard deviations derived from these experiments.
Focusing on an easily interpretable measure, the accuracy means, we observe a variability in performance across LLMs and tasks and identify three main groups of LLMs:
[…]
  • Accuracy means $\geq 80$%: Qwen 2.5 7B, DeepSeek R1, and Devstral 24B fall into this group. The box plot indicates that Qwen 2.5 7B has good capabilities in executing readiness and navigation actions, but produces mixed results when evaluating assertions. DeepSeek R1 is unable to execute navigation actions despite being a model that supports tool calling. It shows good capabilities in evaluating readiness and assertions but it is less effective than Llama 3.1 70B.
  • Accuracy means <80<80%: the remaining four models Qwen 2.5 3B, Mistral 7B, Mistral Nemo 12B, and Qwen 3 14B can execute some NL test cases correctly, but they frequently fail to execute navigation actions. Among them, Qwen 3 14B is close to the threshold, with mean accuracies of 78% for navigation, 98% for readiness, and 84% for assertions. While it succeeds on many navigation tasks, it struggles more than Llama 3 70B, particularly on certain websites (e.g., the personal web site). Qwen 2.5 3B seems to be effective for GUI interaction but performs poorly in evaluating assertions. This reveals a sort of bias in our experiments. We indeed observed that failed assertion evaluations may sometimes come from navigation actions that the LLM agent does not execute correctly. Some LLM agent can indeed return a success in performing a navigation action despite an incorrect real execution. Since the GUI state then diverges from the expected one, the assertions fail. Our current NL test case dataset, TestG, does not allow precise detection of navigation errors. This is why we also use TestA in these experiments.
@@
Setup: In RQ1, we measured accuracy means and standard deviations for several LLM agents. Based on these results, we selected three representative models: Llama 3 70B, as the most capable LLM in executing NL test cases; Mistral Nemo 12B, as the least capable; and Qwen3 14B, which demonstrated mixed performance.
@@
Aggregating across all test suites and LLMs, we obtain a MRE mean of 11%This indicates that, on average, our estimated measure deviates from the observed stability of NL test case executions by a non-negligible margin. However, Figure 7 shows a pronounced difference in MRE between the models. Mistral Nemo 12B shows the highest MRE mean of 30%, whereas Llama 3.1 70B and Qwen 3 14B achieve much lower values, with a MRE mean of 2%.
@@ 6.3 Threats To Validity
1) The $AUT$s we considered in our experiments consist of only 5 different web sites. While we included some negative assertions in the test suites, the $AUT$s themselves are not supposed faulty. To strengthen external validity, further kinds of applications, such as mobile applications, should be considered. 2) We considered 8 LLMs that can be deployed on local servers, but excluded cloud-based LLMs that may provide better results for GUI interaction. We only focused in this paper on LLMs that can be deployed on local servers as they offer the advantage of ensuring confidentiality. These LLMs come from different companies, with model sizes ranging from 3B to 70B parameters. However, future work could explore the evaluation of our tools with further LLMs. 3) We used the framework Stagehand for performing navigation actions, which is relatively new and still evolving. New versions of Stagehand may yield better results, especially for small LLMs (e.g., 7B LLMs). Other frameworks could also be considered. For example, we conducted some experiments with BrowserUse, which is a framework rather dedicated to the execution of general scenarios. However, our initial results were not promising.
@@ 7 Conclusion
In this paper, we investigated the feasibility of executing NL test cases for GUI applications using LLM agents. To support this study, we introduced formal notions of weak unsoundness and proposed the use of Six-Sigma levels to characterise the abilities of LLM agents. In addition, we defined a complementary measure to quantify the consistency of NL test case execution.
Further, we designed several NL test suites and developed prototype tools to execute these test cases. Experiments across 8 LLMS of varying sizes showed that current LLMs can handle GUI testing, under some conditions: no ambiguity in actions and assertions, and LLMs being able to perform navigation and evaluation tasks with high accuracy and low standard deviation. The evaluation also highlighted important limitations. While Llma 3.1 70B showed promising results, smaller LLMs struggled with navigation or assertion evaluation. Our approach has a number of ways it can be improved, including more efficient tools for GUI data extraction and more coverage of NL test case interactions.
This work also opens future directions. Can NL test cases be generated from higher-level scenarios or navigation maps? Can small LLMs be fine-tuned to specialise in some specific tasks, while larger ones only handle complex tasks? If agent compositions could be optimised dynamically to GUI complexity, this might provide more efficient executions. Can screenshots be included in test suites to better detect navigation issues? Can one extend our tools to balance properties like soundness, laxness, controllability, and performance? Can NL test cases be designed to evaluate non-functional aspects, such as security? Can one improve our tools to optimise GUI data extraction, handle response timeouts, ans support more complex assertion structures?

### v2

@@ 1 Introduction
In this paper, we propose an algorithm and contexts under which NL test case unsoundness becomes acceptable. We formulate this notion of acceptability and define measures to evaluate test case execution consistency. Specifically, we propose an algorithm for executing NL test cases with guardrail mechanisms. It dynamically completes the test steps and controls them, enabling evaluation of whether each step is executed as intended. We leverage a set of specialised agents to handle specific actions: navigation through the GUI, readiness of the GUI to perform the next step, and evaluation of assertions.
To assess the capabilities of these agents in performing these actions, we define three measures. In addition, we introduce another measure to quantify the execution consistency of NL test cases—that is, the ability of our algorithm to produce stable and repeatable test outcomes.
@@
  • two prototype tools and corresponding test suites, made publicly available in [Salva and Taguelmimt, 2025b].
@@ 3 Test Case Execution
With this in mind, we propose an algorithm that supplements on the fly a NL test case with additional internal guardrail actions. These actions aim at orchestrating the execution of agents and supervise how the $AUT$ behaves, ensuring that every navigation action is both feasible and carried out as expected. Specifically, we introduce two internal actions, denoted readinessreadiness and observeobserve, which are systematically injected for every navigation action, as illustrated in Figure 4(a). These actions are said to be internal in the sense that they are not observable during execution. readinessreadiness checks whether the next navigation action is executable, i.e., whether the necessary UI elements are present on the interface to perform the action. In turn, observeobserve checks whether the last executed navigation action has led to a visible change in the GUI and whether the agent has successfully completed it. Both internal actions also contribute to evaluating the consistency and reliability of the test case execution (Section 4). In addition to these internal actions, our algorithm adds to the test case the notion of inputs, i.e. navigation actions, starting by the symbol ”?” applied or sent to $AUT$, and outputs, i.e. GUIs initiated by $AUT$, starting by the symbol ”!” (see Figure 4(a)).
@@ 15   Add $q_{i-1}\xrightarrow{error}INC$ to $\rightarrow$;
All this is reflected in Algorithm 1. It incrementally unfolds the NL test case $tc$ into a tree structure expressing the synchronous execution of $tc$ on $AUT$. We propose to model this execution by using an Input Output Labelled Transition System (IOLTS) denoted $tc|AUT$ whose edges are labelled with actions and terminal nodes are labelled with test verdicts, $tc|AUT=<Q,L\cup\mathcal{T},\rightarrow,q_{0}>$, with $Q$ the state set, $L=L_{I}\cup L_{O}$ the set of inputs and outputs, respectively, $\mathcal{T}$ the set of internal actions, $\rightarrow$ the transition set and $q_{0}$ the initial state.
Algorithm 1 starts by covering every navigation action $a_{i}$ in $a_{1}\dots a_{k}$. It encapsulates the action $a_{i}$ flanked by its corresponding actions $readiness(a_{i})$ and observeobserve, as illustrated in Figure 4(a). If either $readiness(a_{i})$ or observeobserve are evaluated to false, the test case cannot be executed because the action cannot be performed or has not updated the GUI, and an Inconclusive verdict is returned. Otherwise, the algorithm continues with the next action. When the algorithm reaches the assertions $\mathcal{A}_{k+1}\dots\mathcal{A}_{l}$, it incrementally evaluates them, as illustrated in Figure 4(b). If an assertion evaluates to false, then a Fail verdict is returned. The Inconclusive verdict is also reached in case an error is triggered by an agent. Otherwise, a Pass verdict is given.
@@ 3.1 Navigation Action Execution
The agent is invoked using prompts we wrote and structured following the patterns proposed in [White et al., 2023]. We here considered: 1. the “Fact Checklist” pattern that asks the agent to output a list of facts included in the final answer, thereby enhancing transparency on the results provided; 2. the “Template” pattern that enforces an output format in JSON, which includes fields such as extracted facts, results, and a boolean indicator of task completion status; 3. The “Recipe” pattern, which builds a complete sequence of steps given some partially “ingredients” to achieve the given task. The prompt is given in Annex in Figure 8.
@@ 3.3 Assertion Evaluation
Table 2: Examples of assertions expressed with formulas.
@@ 4 Test Case Execution Consistency
We recall that a test case is made up of four kinds of actions: readiness, navigation, assertion, and observation. Among these, three are performed by agents—readinessreadiness, navigation actions, and assertions. To measure the consistency of their executions, we define three measures as follows:
First, $s_{r}(a)$ measures the consistency of $readiness(a)$, with $a$ a navigation action. Recall that the action $readiness(a)$ is either performed by readiness_strict(a)readiness\_strict(a) or readiness_w_agent(a)readiness\_w\_agent(a). readinessreadiness _strict\_strict is expressed by a formula and is consistent. When the formula evaluates to false, we use readiness_w_agent(a)readiness\_w\_agent(a), which itself calls $agent_{readiness}$. The execution consistency of readiness_wreadiness\_w _agent(a)\_agent(a) is related to the standard deviation of $agent_{readiness}$. $s_{r}$ is hence defined as:
$0\leq s_{r}(a)\leq 1=_{def}\left\{\begin{array}[]{l}1\text{ if }readiness\_strict(a)=1\\ 1-2\sigma(agent_{readiness})\text{ otherwise }\end{array}\right$.
@@
Third, $s_{a}(\mathcal{A})$ measures the consistency of the evaluation of $\mathcal{A}$. Similarly to readinessreadiness, this action is either performed by assertassert $strict(\mathcal{A})$ or $sert\_w\_agent(\mathcal{A}). sas_{a}$ is hence defined as:
$0\leq s_{a}(\mathcal{A})\leq 1=_{def}\left\{\begin{array}[]{l}1\text{ if }assert\_strict(\mathcal{A})=1\\ 1-2\sigma(agent_{assert})\text{ otherwise }\end{array}\right$.
@@ 5 Test Case Unsoundness
Let $tc=a_{1}\dots a_{k}\mathcal{A}_{k+1}\dots\mathcal{A}_{l}$ be a NL test case such that there exists $a_{1}!g_{1}\dots?a_{k}!g_{k}\in traces(S)$, and $\mathcal{A}_{j}(k+1\leq j\leq l)$ are true on $g_{k}$, and $tc$ is made up of atomic actions and assertions. If $(\sigma(agent_{nav})<3\sigma$, $\sigma(agent_{eval})<3\sigma$, $\sigma(agent_{assert})<3\sigma)\in\Omega_{r}$, then $tc$ is weakly unsound.
@@
Proofs of these propositions are given in [Salva and Taguelmimt, 2025a].
@@ 6 Experimental Results
To perform these experiments, we developed 2 prototype tools composed of three agents. Two agents aim at evaluating readiness actions and assertions, as described in Section 3. The last agent was built on top of the Stagehand  [^3] https://www.stagehand.dev/ framework, which automates browser interactions with natural language and code by means of LLMs. Stagehand was used to perform navigation actions and, in part, for data extraction from pages:
@@
  • NLTestRunner implements Algorithm 1. Given a test suite and a number of runs $N$, it executes NL test cases, returns verdicts, and computes the estimated consistency measures and the observed test case execution consistencies if $N>0$. It can recognise 12 prompt forms to perform strict readiness and strict assertion evaluations.
We performed our experiments with 8 LLMs, whose sizes range between 3B and 70B parameters. The LLM inference was carried out using Ollama [^4] https://ollama.com/, installed on a local server equipped with an NVIDIA L40 GPU (48 GB memory). All Web sites, test suites, and tools are available in [Salva and Taguelmimt, 2025b].
@@
Setup: To address this question, we considered these 8 LLMs: Qwen2.5:3B, Qwen2.5 7B, Mistral 7B, Mistral Nemo 12B, Deepseek-R1 14B, Qwen3:14B, Mistral Devstral 24B, and Llama3.1:70B. These models, provided by different companies, range from 3B to 70B parameters. We selected LLMs that support tool calling, which is a required feature to allow GUI interaction.
@@
Results: Figures 5 and 6 present box-plots of the distributions of mean accuracy across the NL test cases. The plots also show quartiles and mean values (indicated with crosses) of performing navigation actions, returning expected readiness, and returning expected assertion verdicts. Table 3 lists the standard deviations derived from these experiments.
Focusing on an easily interpretable measure, the mean accuracy, we observe a variability in performance across LLMs and tasks and identify three main groups of LLMs:
  • Mean accuracies >93.32>93.32% (level $3\sigma$): considering all test case step categories, only one model demonstrated acceptable capabilities in running NL test cases. Llama 3.1 70B achieved the best performance in our experiments, with mean accuracies greater than or equal to 98%. However, the box plot reveals some outliers in assertion evaluations;
  • Mean accuracies $\geq 80$%: Qwen 2.5 7B, DeepSeek R1, and Devstral 24B fall into this group. The box plot indicates that Qwen 2.5 7B has good capabilities in executing readiness and navigation actions, but produces mixed results when evaluating assertions. DeepSeek R1 is unable to execute navigation actions despite being a model that supports tool calling. It shows good capabilities in evaluating readiness and assertions but it is less effective than Llama 3.1 70B.
  • Mean Accuracies <80<80%: the remaining four models Qwen 2.5 3B, Mistral 7B, Mistral Nemo 12B, and Qwen 3 14B can execute some NL test cases correctly, but they frequently fail to execute navigation actions. Among them, Qwen 3 14B is close to the threshold, with a mean accuracy of 78% for navigation, 98% for readiness, and 84% for assertions. While it succeeds on many navigation tasks, it struggles more than Llama 3 70B, particularly on certain websites (e.g., the personal website). Qwen 2.5 3B seems to be effective for GUI interaction but performs poorly in evaluating assertions. This reveals a sort of bias in our experiments. We indeed observed that failed assertion evaluations may sometimes come from navigation actions that the LLM agent does not execute correctly. Some LLM agent can indeed return a success in performing a navigation action despite an incorrect real execution. Since the GUI state then diverges from the expected one, the assertions fail. Our current NL test case dataset, TestG, does not allow precise detection of navigation errors. This is why we also use TestA in these experiments.
@@
Setup: In RQ1, we measured the mean accuracies and standard deviations for several LLM agents. Based on these results, we selected three representative models: Llama 3 70B, as the most capable LLM in executing NL test cases; Mistral Nemo 12B, as the least capable; and Qwen3 14B, which demonstrated mixed performance.
@@
Aggregating across all test suites and LLMs, we obtain a mean MRE of 11%. This indicates that, on average, our estimated measure deviates from the observed stability of NL test case executions by a non-negligible margin. However, Figure 7 shows a pronounced difference in MRE between the models. Mistral Nemo 12B shows the highest mean MRE of 30%, whereas Llama 3.1 70B and Qwen 3 14B achieve much lower values, with a mean MRE of 2%.
@@ 6.3 Threats To Validity
1) The $AUT$s we considered in our experiments consist of only 5 different web sites. Furthermore, while we included some negative assertions in the test suites, the $AUT$s themselves are not supposed faulty. To strengthen external validity, further kinds of applications, such as mobile applications, should be considered. 2) We considered 8 LLMs that can be deployed on local servers, but excluded cloud-based LLMs that may provide better results for GUI interaction. We only focused in this paper on LLMs that can be deployed on local servers as they offer the advantage of ensuring confidentiality. These LLMs come from different companies, with model sizes ranging from 3B to 70B parameters. However, future work could explore the evaluation of our tools with further LLMs. 3) We used the framework Stagehand for performing navigation actions, which is relatively new and still evolving. New versions of Stagehand may yield better results, especially for small LLMs (e.g., 7B LLMs). Other frameworks could also be considered. For example, we conducted some experiments with BrowserUse, which is a framework rather dedicated to the execution of general scenarios. However, our initial results were not promising.
@@ 7 Conclusion
In this paper, we investigated the feasibility of executing NL test cases for GUI applications using LLM agents. To support this study, we introduced formal notions of weak unsoundness and proposed the use of Six-Sigma levels to characterise the abilities of LLM agents. In addition, we defined a complementary measure to quantify the consistency of NL test case execution. Building on this, we developed an algorithm with guardrail mechanisms and specialised agents that dynamically verify the test step executions. To evaluate our approach, we designed several NL test suites and implemented prototype tools to execute them. Experiments across 8 LLMS of varying sizes showed that current LLMs can handle GUI testing, under some conditions: no ambiguity in actions and assertions, and LLMs being able to perform navigation and evaluation tasks with high accuracy. As we demonstrated, Meta Llama 3.1 70B surpasses the 3-Sigma threshold in NL test case execution with a high level of execution consistency, while smaller LLMs struggled with navigation or assertion evaluation.
This work also opens several future directions. Can NL test cases be generated from higher-level scenarios or navigation maps? How can our tools be improved to optimise GUI data extraction, handle response timeouts, and support more complex assertion structures? Can screenshots be included into test sets to better evaluate agent capabilities? How can our algorithm be extended to balance properties like soundness, laxness, controllability, and efficiency? Can NL test cases be designed to evaluate non-functional aspects, such as security? Can small LLMs be fine-tuned to specialise in some specific tasks, while larger ones only handle complex tasks? If agent compositions could be optimised dynamically to GUI complexity, this might provide more efficient executions.